Show that any nilpotent element is quasi-regular in every ring.
1
Expert's answer
2012-10-30T10:26:52-0400
Say an+1= 0. Then
a ◦ (−a − a2 −· · ·−an)= −a2 −· · ·−an + a(a + a2+ · · · + an) = 0,
and similarly
(−a − a2 −· ··−an) ◦ a = 0.
So, element a is quasi-regular.